OpenStudy (anonymous):

A cylindrical grain bin has a radius of 9 feet and a height of 30 feet. What is the surface area of the bin?

OpenStudy (anonymous):

the surface area of the cylinder top is pi*r^2 the diameter of the circle is 2*9=18, so the cylinder wall is just b*h = (18*pi) * 30 Then add the surfaces. Usually you will use double the circle area since a cylinder has a top and bottom, but it's a silo on the ground, so i don't know if that would be included